02[height = 2cm]dc5logodebconf5 − 1logo

Updating scanners and filters in Debian stable (3.1)
a separate volatile section
volatile.debian.org

1

security.debian.org
ftp.debian.org
volatile.debian.net
deb http://volatile.debian.org/debian-volatile sarge/volatile main
deb-src http://volatile.debian.org/debian-volatile sarge/volatile main
deb http://volatile.debian.org/debian-volatile sarge/volatile-sloopy main
deb-src http://volatile.debian.org/debian-volatile sarge/volatile-sloopy main
http://volatile.debian.net
debian-volatile@lists.debian.org
irc.oftc.net - #debian-volatile

Joeyhttp://lists.debian.org/debian-devel/2004/09/msg00795.html http://lists.debian.org/debian-devel/2004/09/msg00854.html http://lists.debian.org/debian-devel/2004/09/msg01304.html

Bug Tracking System